Circular No. 7872

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

XTE J1751–305 M. Ehle, M. G. Breitfellner, L. Tomas, M. Santos Lleo and E. Ver-

dugo, XMM-Newton Science Operations Centre, European Space Agency, Madrid, report: “XMM-Newton observed the location of the recently discovered millisecond pulsar XTE J1751–305 (IAUC 7867) during Apr. 7.526–7.938 UT. The details of the instrument setup are provided at

http://xmm.vilspa.esa.es/obs-public/obs view.tcl (revolution 426).

Preliminary analysis of the EPIC-MOS2 gives a source position of α = 17^{h}51^{m}13^{s}.5, δ = −30^{o}37′22′′ (equinox 2000.0). At this stage of data reduction the position error is expected to be < 10′′. The approximate 0.5–12keV x-ray countrate measured with EPIC-PN was 61.3 ± 0.1 counts/s. No systematic flux trends were observed, and no x-ray bursts were detected. Analysis of the EPIC-MOS1 and EPIC-PN data obtained in timing mode to investigate the pulse period of XTE J1751–305 is underway. The observation data files and pipeline products will be made public immediately

after finalization at http://xmm.vilspa.esa.es.”

P. G. Jonker and G. Nelemans, Cambridge University; P. Groot, University of Nijmegen; M. van der Klis, University of Amsterdam; J. Hjorth, University of Copenhagen; A. Delsanti, European Southern Observatory; and B. Stappers, ASTRON, report: “On comparing a 15-min R-band image obtained with the 1.54-m Danish telescope at La Silla on Apr. 8.37 UT with a DSS2 (1991 Aug. 31) R-band plate image, we found an object that has brightened by > 1 mag in the error circle of the millisecond x-ray pulsar XTE J1751–305 (IAUC 7867, 7870). Astrometry using USNO-A2.0 astrometric standards yielded a preliminary position of α = 17^{h}51^{m}16^{s}.1, δ = −30^{o}38′00′′ (equinox 2000.0, 2′′ uncertainty). Differential photometry yields a current estimate of R = 17.5.”

In a subsequent report, Jonker et al. add: “A spectrum obtained by M. Calkins on Apr. 9.50 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph) of the variable mentioned above shows it to be a late-type M star. Therefore, it is most likely not the counterpart of XTE J1751–305.”

SUPERNOVA 2002bw IN PGC 59990 T. Matheson, S. Jha, P. Challis and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002bw (cf. IAUC 7864), obtained by Calkins as above on Apr. 4.49 UT, shows it to be a type-Ia supernova ∼ 3 weeks after maximum.

2002 April 9  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Brian G. Marsden